In order to complete an experiment in chemistry lab, Sylvia needs to make a mixture of 130 mL of a 37% base solution from a 67% base solution and a 24% base solution. Find the volume of the 67% base solution that Sylvia needs to use to make the mixture.
Find the volume of the 67% Base solution.

Answer by ankor@dixie-net.com(22740)

In order to complete an experiment in chemistry lab, Sylvia needs to make a mixture of 130 mL of a 37% base solution from a 67% base solution and a 24% base solution.
Find the volume of the 67% base solution that Sylvia needs to use to make the mixture.
Find the volume of the 67% Base solution.
:
let x = amt of 67% solution
The resulting amt is to be 130 ml, therefore
(130-x) = amt of 24% solution
A typical mixture equation
.67x + .24(130-x) = .37(130)
.67x + 31.2 - .24x = 48.1
.67x - .24x = 48.1 - 31.2
.43x = 16.9
x = 16.9/.43
x = 39.3 ml of the 67% solution required
then
130 - 39.3 = 90.7 of the 24% solution required